# Counting points on character varieties

### v1

@@ Abstract.
We count points on the character varieties associated with punctured surfaces and regular semisimple generic conjugacy classes in reductive groups. We find that the number of points are palindromic polynomials. This suggests a $P=W$ conjecture for these varieties. We also count points on the corresponding additive character varieties and find that the number of points are also polynomials, which we conjecture have non-negative coefficients. These polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers.
@@ 1.1. Overview
Character varieties of surface groups play a central role in diverse areas of mathematics and physics, such as gauge theory [AtiyahBott83, HitchinProceedingofLMS87], mirror symmetry [HauselThaddeusInventiones2003], non-abelian Hodge theory [Simpson92, Simpson94i], and the geometric Langlands program [BeilinsonDrinfeld, BenZviNadler]. Consequently, the topology and geometry of these varieties have been the focus of active research for decades. In their groundbreaking works [HRV08, HLRV11], Hausel, Letellier, and Rodriguez-Villegas counted points on $\mathrm{GL}_{n}$-character varieties, leading to deep conjectures about the geometry of these varieties. Many of these conjectures have since been proven, greatly elucidating the geometry of character varieties. Notable achievements include the determination of the Betti numbers [SchiffmannAnnals2016, MellitInventiones21, Mellit20], the establishment of curious Poincaré duality and curious Hard Lefschetz [MellitDecomposition], and the recent resolution of the $P=W$ conjecture, cf. [HoskinPW].
@@
In [HLRV11], the authors also considered the additive version of character varieties, where the natural commutation relation is replaced by its Lie algebraic analogue. According to the Riemann–Hilbert correspondence, points on character varieties correspond to flat connections on vector bundles. When $g=0$, the points on the additive character variety correspond to flat connections on the trivial bundle. Thus, when $g=0$, the additive character variety is a dense open subvariety of the usual character variety. This direct relationship is lost in higher genus; nevertheless, it was conjectured in [HLRV11] that the cohomology of the additive character variety captures the pure part of the cohomology of the multiplicative character variety in all genus. As evidence, the authors proved that additive character varieties are pure. This result relied on the beautiful bridge built by Crawley-Boevey between additive character varieties and quiver varieties associated to comet-shaped quivers [CBIndecomposable]. Under this correspondence, the number of points on the additive character variety matches the Kac polynomials of the corresponding quiver. Utilising these ideas, Hausel, Letellier, and Rodriguez-Villegas proved Kac’s conjecture for arbitrary quivers [HLVKacAnnals2013].
@@
Despite significant advances in understanding the geometry of character varieties associated with $\mathrm{GL}_{n}$ (and more generally, groups of type $A$), there is limited understanding of character varieties associated with arbitrary reductive groups $G$. Understanding these general cases is crucial for applications in fields such as Langlands duality and mirror symmetry, where the Langlands dual group plays a central role. In this paper, we count points on both multiplicative and additive character varieties associated with punctured surface groups and semisimple conjugacy classes in arbitrary reductive groups. Our work builds upon and generalises the contributions of Cambò [Cambo17], who studied $\mathrm{Sp}_{2n}$-character varieties; [BK22], who investigated character stacks associated with compact surface groups; and [KNP23], who examined character varieties in presence of regular semisimple and regular unipotent conjugacy classes at the punctures.
@@
One of our main results is that the number of points of multiplicative character varieties over finite fields are palindromic polynomials. This hints at a $P=W$ conjecture, a curious Poincaré duality, and a curious Hard Lefschetz property for these varieties. [^1] To make further progress in this direction, one needs a reductive analogue of Markman’s result [MarkmanGenerators2001] that the tautological classes (along with those from the underlying surface) generate the cohomology. Another important contribution of this paper is an explicit polynomial formula for the number of points of additive character varieties. The resulting polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers. Computer calculations suggest that these polynomials have non-negative coefficients [^2] See [Bailey] for the computer code written using CHEVIE.; however, lack of direct connection with quiver varieties makes proving this a nontrivial task. Finally, one of the main contributions of [HRV08, HLRV11] was to relate the number of points of $\mathrm{GL}_{n}$-character varieties with Macdonald polynomials and use this to provide a conjectural formula for the mixed Hodge polynomials of these varieties. [^3] As far as we know, these conjectures (as well as the purity conjecture mentioned above) are still open. Formulating analogous conjectures in the reductive setting is an interesting open problem.
1.1.1. Structure of the text
In the rest of the introduction, we formulate our main results. In §2, we gather some facts about reductive groups and root systems. In §3, we discuss basic results about the geometry of generic character varieties. The main computation regarding points on character varieties appears in §4 while its additive analogue appears in §5. Finally, in §6, we give explicit examples of counting polynomials for groups of low rank and discuss some of the consequences of these formulas.
1.1.2. Acknowledgement
We would like to thank Ivan Cheltsov, Anand Deopurkar, Jack Hall, Ian Le, Emmanuel Letellier, Paul Levy, Jean Michel, and Behrouz Taji for helpful conversations. MK was supported by an Australian Research Council (ARC) discovery grant. GH was supported by the National Research Foundation of Korea (NRF). BW and SG were supported by Australian postgraduate scholarships. We would like to thank the Sydney Mathematical Research Institute for facilitating a visit by Emmanuel Letellier, giving further impetus to this project. The material of §4 (resp. §5) forms a part of the Master’s thesis of BW (resp. SG).
@@ 1.2. Generic character varieties
$\Gamma=\frac{\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\rangle}{[a_{1},b_{1}]\cdots[a_{g},b_{g}]c_{1}\cdots c_{n}}.$
@@
This is an affine scheme of finite type over $\mathbb{F}_{q}$ with an action of $G$ by conjugation. The associated character variety is
@@
  1. (1)
    $\prod_{i=1}^{n}C_{i}\subseteq[G,G]$ (otherwise, $R$ would be empty).
  2. (2)
    Each conjugacy class $C_{i}$ contains an element $S_{i}\in T(\mathbb{F}_{q})$ which is strongly regular; i.e., $C_{G}(S_{i})=T$.
  3. (3)
@@
We also assume $2g+n\geq 3$, otherwise the generic assumption implies that $X$ is empty. As explained in §3.2, the generic assumption ensures that the action of $G/Z$ on $R$ has finite étale stabilisers and that $R$ is smooth and equidimensional. Thus, ${X}$ is also equidimensional of pure dimension
$\dim({X})=(2g-2+n)\dim(G)+2\dim(Z)-n.\dim(T).$
We assume $Z=Z(G)$ is connected, the ground characteristic is very good for $G$ and $q$ is large enough (see Convention 25). Then, we have:
@@
$|X(\mathbb{F}_{q})|$ is a polynomial in $q$ given explicitly in §4.7. This is the $E$-polynomial [^4] See, e.g., [LetellierVillegas] for recollections on $E$-polynomials of varieties over $\mathbb{F}_{q}$. of $X$ and depends only on $(G,g,n)$; i.e., it is independent of the conjugacy classes $C_{i}$ (provided they satisfy Assumption 1). Moreover, this polynomial is palindromic.
The starting point of the proof is the observation that since $X$ is the coarse moduli space of the Deligne-Mumford stack $[X]$, we have $|X(\mathbb{F}_{q})|=|[X](\mathbb{F}_{q})|$. An application of Lang’s theorem then implies
@@
Next, the Frobenius Mass Formula allows one to write $|R(\mathbb{F}_{q})|$ in terms of irreducible complex representations of the finite group $G(\mathbb{F}_{q})$. We then use results of Deligne and Lusztig to analyse this formula and prove the theorem; see §4 for details.
@@
    We expect that the theorem holds without assuming that $Z(G)$ is connected; however, the expression for $|X(\mathbb{F}_{q})|$ becomes more complicated due to disconnected centralisers in the dual group.
@@
    Our analysis shows that $|[{X}](\mathbb{F}_{q})|$ is a polynomial in $q$ even if we drop the generic assumption. We expect that the regularity assumption is also not required for this. We emphasise, however, that without the generic assumption, the variety $X$ is quite different from the stack $[X]$.
@@
    The leading coefficient of the $E$-polynomial equals the number of connected components of ${X}$. It is easy to analyse this polynomial to conclude that $|\pi_{0}({X})|=|\pi_{0}(Z(\check{G}))|$, where $\check{G}$ is the Langlands dual group of $G$; see [Whitbread24, §7.5] for details.
@@
    The value at $1$ of the $E$-polynomial gives the Euler characteristic of $X$. This can be used to show that if $g>1$, or $g=1$ and $Z$ is nontrivial, then the Euler characteristic of ${X}$ is 00. Expressions for the Euler characteristic when $g=0$ or $g=1$ and $Z$ is trivial are given in §6.1. For instance, when $G=G_{2}$ and $g=0$, we find
    $\chi(X)=2^{2n-7}3^{n-3}(n-1)(n-2)(207n^{2}-621n+350).$
In lieu of the above theorem, it is reasonable to expect:
@@
We expect that the above theorem and conjecture hold without the regularity assumption. However, the generic assumption and semisimplicity are crucial. When $G=\mathrm{GL}_{n}$, Mellit constructed a cell decomposition for ${X}$, thus providing a conceptual understanding of the fact that ${X}$ is polynomial count and proving in addition that it is Hodge–Tate and satisfies curious Hard Lefschetz and curious Poincaré duality [Mellit20]. We expect that ${X}$ has an analogous cell decomposition for arbitrary reductive group $G$.
1.3. Additive analogue
Let $\mathfrak{g}$ be the Lie algebra of $G$ and ${\mathcal{O}=(\mathcal{O}_{1},\ldots,\mathcal{O}_{n})}$ a tuple of $G$-orbits on $\mathfrak{g}$. Define the additive representation variety by
We have the adjoint action of $G$ on ${A}$. We call the resulting GIT quotient
${Y}:={A}/\!\!/G$
the additive character variety associated to $(\Gamma,G,\mathcal{O})$. We will be working under the obvious additive analogues of Assumption 1. Then the action of $G/Z$ on ${A}$ has finite étale stabilisers, ${A}$ is smooth and equidimensional, and ${Y}$ is also equidimensional of the same dimension as ${X}$. As above, we assume that $Z$ is connected, the ground characteristic is very good and $q$ is large enough (see Convention 26). Then we have:
Theorem 5.
$|Y(\mathbb{F}_{q})|$ is a polynomial in $q$ given explicitly in §5.6. This is the $E$-polynomial of $Y$ and depends only on $(G,g,n)$; i.e., it is independent of the orbits $\mathcal{O}_{i}$ (provided they satisfy the additive analogue of Assumption 1). Moreover, this polynomial is monic; thus, $X$ is connected.
The starting point of the proof is an additive analogue of the Frobenius formula which gives an expression for $|{Y}(\mathbb{F}_{q})|$ in terms of the Fourier transforms of $G$-invariant orbits on $\mathfrak{g}$. We then use a result of Kazhdan and Letellier to evaluate these Fourier transforms and compute $|Y(\mathbb{F}_{q})|$; see §5 for details. Using the formula for $|Y(\mathbb{F}_{q})|$, we can compute the Euler characteristic $Y$. For instance, for $G=G_{2}$, we find that $\chi(Y)$ equals
$\frac{1}{6}(6-2^{n}+19\cdot 3^{n+1}+5^{n})+395\cdot 12^{n-2}+5g\cdot 6^{n}+6^{n-2}(48-65\cdot 2^{n})n+5\cdot 3^{n}4^{n-2}(16g^{2}+8g(n-3)+n^{2}).$
When $G=\mathrm{GL}_{n}$, the additive character variety $Y$ is isomorphic to a quiver variety and $|Y(\mathbb{F}_{q})|$ is a polynomial with non-negative coefficients [HLRV11]. Computer calculations suggest that this fact generalises to the reductive setting:
Conjecture 6.
The variety ${Y}$ is pure; thus, $|{Y}(\mathbb{F}_{q})|$ has non-negative coefficients.
As noted in the overview section, when $G=\mathrm{GL}_{n}$, $|{Y}(\mathbb{F}_{q})|$ is the Kac polynomial of a certain comet shaped quiver. The coefficients of Kac polynomials are expected to be the graded dimensions of the associated graded Borcherd algebra [Schiffmann18, §3]. It would be interesting to find a similar interpretation for the coefficients of $|{Y}(\mathbb{F}_{q})|$ for general reductive groups.
2. Root systems and reductive groups
In this section, we recall basic facts and notations about root systems and reductive groups. The material here is standard, though perhaps the facts about pseudo-Levi subgroups and quasi-isolated semisimple elements are not as widely known.

### v2

# Counting points on generic character varieties

@@ Abstract.
We count points on character varieties associated with punctured surfaces and regular semisimple generic conjugacy classes in reductive groups. We find that the number of points are palindromic polynomials. This suggests a $P=W$ conjecture for these varieties. We also count points on the corresponding additive character varieties and find that the number of points are also polynomials, which we conjecture have non-negative coefficients. These polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers.
@@ 1.1. Overview
Character varieties of surface groups play a central role in diverse areas of mathematics and physics, such as gauge theory [AtiyahBott83, HitchinProceedingofLMS87], mirror symmetry [HauselThaddeusInventiones2003], non-abelian Hodge theory [Simpson92, Simpson94i], and the geometric Langlands program [BeilinsonDrinfeld, BenZviNadler]. Consequently, the topology and geometry of these varieties have been the focus of active research for decades. In their groundbreaking works [HRV08, HLRV11], Hausel, Letellier, and Rodriguez-Villegas counted points on $\mathrm{GL}_{n}$-character varieties, leading to deep conjectures about the geometry of these varieties.
Many of these conjectures have since been proven. Notable achievements include the determination of the Betti numbers [SchiffmannAnnals2016, MellitInventiones21, Mellit20], the establishment of curious Poincaré duality and curious Hard Lefschetz [MellitDecomposition], and the recent resolution of the $P=W$ conjecture [hausel2022p, maulik, HoskinPW].
@@
In [HLRV11], the authors also considered the additive version of character varieties, where the natural commutation relation is replaced by its Lie algebraic analogue. According to the Riemann–Hilbert correspondence, points on the usual (multiplicative) character varieties correspond to flat connections on vector bundles. When $g=0$, the points on the additive character variety correspond to flat connections on the trivial bundle. Thus, when $g=0$, the additive character variety is a dense open subvariety of the usual character variety, sometimes known as the open de Rham space [HWW23].
This direct relationship between additive and multiplicative character varieties is lost in higher genus; nevertheless, it was conjectured in [HLRV11] that the cohomology of the additive character variety captures the pure part of the cohomology of the multiplicative character variety in all genus. As evidence, the authors proved that additive character varieties are pure. This result relied on the beautiful bridge built by Crawley-Boevey between additive character varieties and quiver varieties associated to comet-shaped quivers [CBIndecomposable]. Under this correspondence, the number of points on the additive character variety matches the Kac polynomials of the corresponding quiver. Utilising these ideas, Hausel, Letellier, and Rodriguez-Villegas proved Kac’s conjecture for arbitrary quivers [HLVKacAnnals2013].
@@
Despite significant advances in understanding the geometry of character varieties associated with $\mathrm{GL}_{n}$ (and more generally, groups of type $A$), there is limited understanding of character varieties associated with arbitrary reductive groups $G$. Understanding these general cases is crucial for applications in fields such as Langlands duality and mirror symmetry, where the Langlands dual group plays a central role.
In this paper, we count points on both multiplicative and additive character varieties associated with punctured surface groups and regular semisimple conjugacy classes in arbitrary reductive groups. Our work builds upon and generalises the contributions of Cambò [Cambo17], who studied $\mathrm{Sp}_{2n}$-character varieties; [BK22], who investigated character stacks associated with compact surface groups; and [KNP23], who examined character varieties in presence of regular unipotent conjugacy classes at the punctures.
@@
One of our main results is that the number of points of multiplicative character varieties over finite fields are palindromic polynomials. This hints at a $P=W$ conjecture, a curious Poincaré duality, and a curious Hard Lefschetz property for these varieties. [^1] To make further progress in this direction, one needs a reductive analogue of Markman’s result [MarkmanGenerators2001] that the tautological classes (along with those coming from the underlying surface) generate the cohomology. Another important contribution of this paper is an explicit polynomial formula for the number of points of additive character varieties. The resulting polynomials can be considered as the reductive analogues of the Kac polynomials of comet-shaped quivers. Computer calculations suggest that these polynomials have non-negative coefficients [^2] See [Bailey] for the computer code written using CHEVIE [GHLMP96].; however, lack of direct connection with quiver varieties makes proving this a nontrivial task.
Finally, one of the main contributions of [HRV08, HLRV11] was to relate the number of points of $\mathrm{GL}_{n}$-character varieties with Macdonald polynomials and use this to provide a conjectural formula for the mixed Hodge polynomials of these varieties. [^3] As far as we know, these conjectures (as well as the purity conjecture mentioned above) are still open. Formulating analogous conjectures in the reductive setting is an interesting open problem.
@@ 1.2. Generic character varieties
$\Gamma={\langle a_{1},b_{1},\ldots,a_{g},b_{g},c_{1},\ldots,c_{n}\,|\,{[a_{1},b_{1}]\cdots[a_{g},b_{g}]c_{1}\cdots c_{n}}=1\rangle}.$
@@
This is an affine scheme of finite type over $\mathbb{F}_{q}$ with an action of $G$ by conjugation. The associated character variety is the GIT quotient
@@
  • The centre $Z=Z(G)$ is connected and the ground characteristic is very good for $G$.
  • $\prod_{i=1}^{n}C_{i}\subseteq[G,G]$, otherwise $R$ would be empty.
  • Each conjugacy class $C_{i}$ contains an element $S_{i}\in T(\mathbb{F}_{q})$ which is strongly regular; i.e., $C_{G}(S_{i})=T$ (cf. [Steinberg65, Humphreys95]).
@@
  • $2g+n\geq 3$, otherwise the generic assumption implies that $X$ is empty.
As explained in §3.2, the generic assumption ensures that the action of $G/Z$ on $R$ has finite étale stabilisers and that $R$ is smooth and equidimensional. Thus, $[X]$ is a smooth Deligne–Mumford stack. Moreover, $X$ and $[X]$ are equidimensional of pure dimension
$\dim({X})=(2g-2+n)\dim(G)+2\dim(Z)-n\cdot\dim(T).$
Before stating our main theorem, we need to recall some facts about polynomial count varieties and $E$-polynomials. We refer the reader to [HRV08] or [LetellierVillegas] for further details. A variety $S$ over $\mathbb{F}_{q}$ is said to be polynomial count if there exists a polynomial $f_{S}(t)\in\mathbb{C}[t]$ such that
$|S(\mathbb{F}_{q^{m}})|=f_{S}(q^{m}),\qquad\forall m\in\mathbb{Z}_{>0}.$
In this case, $f_{S}$ is known as the counting polynomial of $S$. An important fact is that if $S$ is polynomial count then its counting polynomial equals the $E$-polynomial of $S$. The latter is a certain specialisation of the mixed Poincaré polynomial of $S$ (defined using étale cohomology). Since the $E$-polynomial has integer coefficients, it follows that the counting polynomial must belong to $\mathbb{Z}[t]$.
The $E$-polynomial gives useful topological information about the variety. For instance, the leading coefficients of the $E$-polynomial equals the number of irreducible components of top dimension, and that the value of the $E$-polynomial at $1$ equals the Euler characteristic.
Our main result states that generic character varieties are, after a finite base change, polynomial count:
@@
There exists a polynomial $f_{X}\in\mathbb{Z}[t]$ (given explicitly in §4.6) and a positive integer $r$ such that
$|X(\mathbb{F}_{q^{rm}})|=f_{X}(q^{rm}),\qquad\forall m\in\mathbb{Z}_{>0}.$
The starting point is the observation that since $X$ is the coarse moduli space of the Deligne-Mumford stack $[X]$, we have $|X(\mathbb{F}_{q})|=|[X](\mathbb{F}_{q})|$. An application of Lang’s theorem then implies
@@
Next, the Frobenius Mass Formula allows one to write $|R(\mathbb{F}_{q})|$ in terms of irreducible complex representations of the finite group $G(\mathbb{F}_{q})$. We then use results of Deligne and Lusztig [DL76] to analyse this formula and prove the theorem; see §4 for details.
Let $\overline{X}$ denote the base change of $X$ to $\overline{\mathbb{F}_{q}}$ and $E(\overline{X})$ the $E$-polynomial of $\overline{X}$. The previous theorem implies that $E(\overline{X})=f_{X}$. By analysing the polynomial $f_{X}$, we will obtain:
Corollary 3.
  1. (a)
    $E(\overline{X})$ is palindromic.
  2. (b)
    $|\pi_{0}(\overline{X})|=|\pi_{0}(Z(\check{G}))|$, where $\check{G}$ is the Langlands dual group.
  3. (c)
    If $g>1$, or $g=1$ and $Z$ is nontrivial, then $\chi(\overline{X})=0$.
Expressions for the Euler characteristic in the case $g=0$ or $g=1$ and $Z=\{1\}$ is given in §5.3. The first part of the above corollary provides evidence for the following:
Conjecture 4.
When $G=\mathrm{GL}_{n}$, Mellit constructed a cell decomposition for ${X}$, thus providing a conceptual explanation of the fact that ${X}$ is polynomial count and proving in addition that it is Hodge–Tate and satisfies curious Hard Lefschetz and curious Poincaré duality [MellitDecomposition]. We expect that ${X}$ has an analogous cell decomposition for arbitrary reductive group $G$.
@@
    The counting polynomial $f_{X}$ depends only on $(G,g,n)$; i.e., it is independent of the conjugacy classes $C_{i}$ (provided they satisfy Assumption 1).
@@
    We expect that the theorem holds without assuming that the centre $Z(G)$ is connected; however, the expression for $|X(\mathbb{F}_{q})|$ becomes more complicated due to disconnected centralisers in the dual group.
@@
    Our analysis shows that the stacky count $|[{X}](\mathbb{F}_{q})|$ is a polynomial in $q$ even if we drop the generic assumption. We expect that the regularity assumption is also not required for the polynomial property. We emphasise, however, that without the generic assumption, the variety $X$ is quite different from the stack $[X]$.
@@
    There are two reasons for the finite base change in the main theorem. First, the inclusion $[G(\mathbb{F}_{q}),G(\mathbb{F}_{q})]\subseteq[G,G](\mathbb{F}_{q})$ may be proper (e.g. for $G=\mathrm{PGL}_{n}$). However, given $S\in[G,G](\mathbb{F}_{q})$, there exists a positive integer $m$, such that $S\in[G(\mathbb{F}_{q^{m}}),G(\mathbb{F}_{q^{m}})]$. Second, if $W^{\prime}$ is a reflection subgroup of $W$, then $T^{W^{\prime}}$ may be disconnected. The group of components $\pi_{0}=\pi_{0}(T^{W^{\prime}})$ is a finite étale group which may be non-split (i.e. the action of Galois group may be nontrivial). This implies that $|\pi_{0}(\mathbb{F}_{q})|$ depends on the residue of $q$ modulo $|\pi_{0}(\overline{\mathbb{F}_{q}})|$. However, as we shall see, the fact that $\mathrm{char}(k)$ is good for $G$ implies that $q$ is co-prime to $|\pi_{0}(\overline{\mathbb{F}_{q}})|$. Thus, there exists a positive integer $n$ such that
    $q^{n}\equiv 1\mod|\pi_{0}(\overline{\mathbb{F}_{q}})|.$
    It follows that $|\pi_{0}(\mathbb{F}_{q^{n}})|=|\pi_{0}(\overline{\mathbb{F}_{q}})|$.
1.3. The additive analogues
We now consider the additive (or Lie-algebraic) analogues of the above discussions. Let $\mathfrak{g}$ be the Lie algebra of $G$ and ${\mathcal{O}=(\mathcal{O}_{1},\ldots,\mathcal{O}_{n})}$ a tuple of adjoint $G$-orbits on $\mathfrak{g}$. Define the additive representation variety by
The group $G$ acts on ${A}$ by the adjoint action. We call the resulting GIT quotient
${Y}:={A}/\!\!/G$
the additive character variety associated to $(\Gamma,G,\mathcal{O})$.
We will be working under the obvious additive analogues of Assumption 1. Then the action of $G/Z$ on ${A}$ has finite étale stabilisers, ${A}$ is smooth and equidimensional, and ${Y}$ is also equidimensional of the same dimension as ${X}$.
Our main result states that the variety $Y$ is, after a quadratic base change, polynomial count:
Theorem 6.
There exists a polynomial $f_{Y}\in\mathbb{Z}[t]$ (given explicitly in §6.6) such that
$|Y(\mathbb{F}_{q^{2m}})|=f_{Y}(q^{2m}),\qquad\forall m\in\mathbb{Z}_{>0}.$
The starting point of the proof is an additive analogue of the Frobenius formula which gives an expression for $|{Y}(\mathbb{F}_{q})|$ in terms of the Fourier transforms of $G$-invariant orbits on $\mathfrak{g}$. We then use a result of Kazhdan and Letellier to evaluate these Fourier transforms and compute $|Y(\mathbb{F}_{q})|$; see §6 for details.
Remark 7.
The reason for the quadratic base change is that computing $|Y(\mathbb{F}_{q})|$ involves the size of a Springer fibre over $\mathbb{F}_{q}$. It was shown by Springer that this is a polynomial if $q$ is large enough. This was refined in [GoodwinRohrle, Theorem 3.10] where it was shown that (under the assumption that $Z$ is connected and $p$ is good for $G$) the size of the Springer fibre is a polynomial in $q$, unless $G$ has a factor of type $E_{8}$, in which case a quadratic base change may be necessary.
Let $\overline{Y}$ be the base change of $Y$ to $\overline{\mathbb{F}_{q}}$. Analysing the counting polynomial $f_{Y}$, we find:
Corollary 8.
  1. (a)
    The variety $\overline{Y}$ is connected.
  2. (b)
    If $G$ is simple of rank less than $5$, $g\leq 10$ and $1\leq n\leq 1000$, then the coefficients of the counting polynomial $f_{Y}$ are non-negative.
Part (b) of this corollary is established with code written using CHEVIE [GHLMP96]. See [Bailey] for the details of this code. The counting polynomial can also be used to give an expression for the Euler characteristic, see §7.4.
When $G=\mathrm{GL}_{n}$, the additive character variety $Y$ is isomorphic to a quiver variety and $|Y(\mathbb{F}_{q})|$ is a polynomial with non-negative coefficients [HLRV11]. Combined with the previous corollary, this provides evidence for the following:
@@ 1.3. The additive analogues
The cohomology $H_{c}^{*}(\overline{{Y}},\mathbb{Q}_{\ell})$ is pure; thus, $|Y(\mathbb{F}_{q})|$ has non-negative coefficients.
If $G=\mathrm{GL}_{n}$ then $Y$ is isomorphic to a quiver variety and the coefficients of $|Y(\mathbb{F}_{q})|$ were recently shown to be the graded dimensions of the associated graded Borcherd algebra [Lucien], proving a conjecture of Schiffmann [Schiffmann18]. It would be interesting to find a similar interpretation for the coefficients of $|{Y}(\mathbb{F}_{q})|$ for general reductive groups.
1.4. Structure of the text
In §2, we gather some facts about reductive groups and root systems. In §3, we discuss basic results about the geometry of generic character varieties. The main computation regarding points on character varieties appears in §4 and its topological implications are discussed in §5. The additive analogues of these results appears in §6. Finally, in §7, we give explicit examples of counting polynomials for groups of low rank.
1.5. Acknowledgement
We would like to thank Ivan Cheltsov, Anand Deopurkar, Jack Hall, Lucien Hennecart, Ian Le, Emmanuel Letellier, Paul Levy, Sebastian Schlegel Mejia, Jean Michel, Travis Schedler, and Behrouz Taji for helpful conversations.
MK was supported by an Australian Research Council (ARC) discovery grant. GH was supported by the National Research Foundation of Korea (NRF). BW and SG were supported by Australian postgraduate scholarships. We would like to thank the Sydney Mathematical Research Institute for facilitating a visit by Emmanuel Letellier, giving further impetus to this project.
The material of §4 (resp. §6) forms a part of the Master’s thesis of BW (resp. SG).
2. Recollections on root systems and reductive groups
In this section, we recall basic facts and notations about root systems and reductive groups. The material here is standard, though perhaps the facts about pseudo-Levi subgroups and quasi-isolated semisimple elements are not as widely known. For these notions, we refer the reader to e.g. [MS03], [Bonnafe05], or [Taylor22].